OpenStudy (anonymous):

what are the solutions of 3x^2-x+7=0 ?

OpenStudy (anonymous):

use quadratic formula..

OpenStudy (anonymous):

can you help me

OpenStudy (campbell_st):

the formula is \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] in your question a = 3, b = -1 and c = 7 so substitute and evalaute... and you'll find the solutions are complex numbers.